# Modeling Issues with Eye Tracking Data

### v4

@@ 4 Modeling
Several authors have noted that the GEE working covariance $V_{i}=A_{i}^{1/2}R_{\alpha}A_{i}^{1/2}$ plays a role analogous to the autocorrelation of latent residual processes in GLM or state-space models for binary data. In particular, [ZegerLiangAlbert1988] demonstrated that serial correlation among binary observations can be represented either by specifying a latent Gaussian process with AR(1) correlation or by imposing the same AR(1) structure on the working correlation matrix in GEE. Related comparisons appear in [HeagertyZeger2000], who show that marginal models in GEE and latent-process conditional models often share identical first- and second-moment structures, though they differ in higher-order behavior and interpretation.
@@
A related model to AR1 is a first-order moving average model (MA). When binary data is composed of runs (streaks of 0s and 1s), it suggests that once a saccade occurs (e.g., a state switch), its influence is immediate but might not persist indefinitely across the entire time series. This seems more natural to a eye-tracking time series than an autogressive process. This model can also be fit with GEE in the forms of an $R$ matrix having a banded Toeplitz structure. I chose a width of 25 which is approximately the median band length for subjects with at least four runs. For this correlation matrix, $\phi_{M}=.95$. The estimating equations are identical to those of the AR1 model. I use this asa point of comparison only. Subjects’ runs vary in number and length, so setting a fixed bandwidth is highly arbitrary.
@@ 6 Survival Modeling
In the fitted model, the baseline hazard is stratified by transition type so that the temporal risk profiles for leaving and entering the target object may differ freely. Predictors such as Privileged and t are allowed to influence each transition differently through interaction terms with the transition indicator. This structure has interpretable effects: the main coefficients describe how covariates affect the hazard of leaving the target, while the interaction coefficients quantify how those effects change during target reacquisition. Main and transition effects and standard errors can be combined to represent total effects. For this analysis, the main focus is on the total effect for the $0\rightarrow 1$ transition. The survival model with RLE-encoded data provides a computationally efficient modeling alternative to AR1 with uncompressed data.
@@ 7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmtoolbox [Vanegas2023] was used for analysis of the AR1/MA25 models, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 model produced positive definite $R$ matrices; however, the determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ and a ridge of 1e-5 (or 1e-2 for MA25) was applied to $R$. For AR1, estimates were nearly identical to those produced in runs where $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was missing by fiat for the last run. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, MA25, and COX models.
@@
Privileged 0.077 0.052 0.052 0.083 0.053
Contrast -0.079 -0.057 -0.041 -0.013 -0.046
Time 0.854 0.813 0.736 0.498
Priv*Time -0.012 0.001 -0.016 0.012
Contrast*Time -0.015 -0.005 0.007 0.028
  • Note.Total effects are reported for the COM model.
@@ 7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and the transition interaction ($y=0\rightarrow y=1)$, and likewise for $Contrast$. It can be seen that the GLM and AR1/MA25 methods provide similar estimates of experimental effects, and the AR1 and COX methods also have similar estimates. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model, and the $Time$ coefficient is strikingly dissimilar to those of the other models. The GLM, AR1, and MA25 models indicate a strong positive relationship between time and the likelihood of being on-target. The variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. For this reason, the COX method did not produce intercepts or time interactions.
It is interesting that the AR1/MA25 models resulted in similar estimates of experimental effects. This is consistent with the the utility of GEE esimtation providing robust estimates given misspecification of the error correlation. Recall that the AR1, MA25, and COX approaches employ variance modeling alternatives to error correlation, while no adjustment for autocorrelation is present in the GLM model.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM provides the lowest, and the AR1/MA25, COX models provide the highest for the experimental variables. The LAG model produces SEs about double of those for the GLM model, but even so, LAG SEs are half the size those for AR1/MA25 and COX. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator. Still, the SEs for the AR1/MA25 and COX models are similar.
@@
Privileged 0.007 0.034 0.036 0.014 0.030
Contrast 0.004 0.021 0.021 0.009 0.018
Time 0.003 0.014 0.027 0.006
Priv*Time 0.007 0.033 0.066 0.015
Contrast*Time 0.005 0.020 0.038 0.009
  • Note. Robust SEs are reported for total effects for the COX model.
7.3 Random Effects and $\phi$
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, MA25, and COX models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were similar. For the AR1, the estimated residual correlation was extremely high. However, this is not unexpected given the median run length of 35 (out of 112) across all subjects. This level of $\phi$ raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient.
Table 4: Random-effects variances and autoregressive parameters.
Model Subject variance Item variance $\phi$ $\alpha$
@@ 7.3 Random Effects and $\phi$
AR1 0.95 (0.948) 0.993
MA25 0.95 (—) 0.967
  • Note. Ridge estimates are reported for AR1 and MA1 with $\phi=.95$. Estimated values of $\phi$ are given in parentheses, but $\phi$ as not estimated for MA25.
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, the AR1/MA25 and COX provide more conservative alternatives for estimating fixed effects. This result raises important and unresolved questions. First, the finding that these two approaches provide similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions.
Of course, the similarity of the AR1/MA25 and COX models does not imply they are correct. Rather it raises the question of how eye-tracking time series can be modeled. The AR1 and moving average models are impractical for longer time series because the determinant of $R$ approaches zero. While a upper boundary can be set manually, this is purely an ad hoc solution. The survival model, or similar models, may provide a way to avoid near singularities due to serial correlation. Moreover, these models are not truly time series models but rather error-correlation models. Based on results from the current study, it is not clear that time series or error correlztion models are required or even suitable for eye-tracking data.
@@
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient seems to have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction the subject variance component as shown in Table 4.
@@
Data pre-processing is another important issue. Eye-tracking systems commonly compress data to avoid creating mammoth output files. Yet even with preliminary machine compression, files may be large. The RLE encoding method for additional compression is convenient for survival analysis and greatly reduces execution time. Pre-processing also includes data cleaning which involves tricky decisions of which experimental records to retain for analysis. This is especially the case in eye-tracking studies because momentary distractions can result in several consecutive seconds of invalid responses.
@@
The results of this study collectively support the existence of experimental effects for the variable $Contrast$. For the AR1, MA25, and COX models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1/MA25 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength, this demonstrates model assumptions can make substantively important differences.
@@ 9 Acknowledgments
I used AI tools in the preparation of this paper. This included preparation of references in biblatex format, review of relevant psychological literature for understanding experimental variables, querying regarding modeling issues, suggesting R code for data file preparation, and obtaining latex code for tables and equations. I also generated three paragraphs of text which were heavily edited prior to finalizing the paper.

### v5

@@ 4 Modeling
Several authors have noted that the GEE working covariance $V_{i}=\alpha A_{i}^{1/2}R_{\phi}A_{i}^{1/2}$ plays a role analogous to the autocorrelation of latent residual processes in GLM or state-space models for binary data. In particular, [ZegerLiangAlbert1988] demonstrated that serial correlation among binary observations can be represented either by specifying a latent Gaussian process with AR(1) correlation or by imposing the same AR(1) structure on the working correlation matrix in GEE. Related comparisons appear in [HeagertyZeger2000], who show that marginal models in GEE and latent-process conditional models often share identical first- and second-moment structures, though they differ in higher-order behavior and interpretation.
@@ 6 Survival Modeling
In the fitted model, the baseline hazard is stratified by transition type so that the temporal risk profiles for leaving and entering the target object may differ freely. Predictors such as Privileged and t are allowed to influence each transition differently through interaction terms with the transition indicator. This structure has interpretable effects: the main coefficients describe how covariates affect the hazard of leaving the target, while the interaction coefficients quantify how those effects change during target reacquisition. For this analysis, the main focus is on the effect for the $0\rightarrow 1$ transition. The survival model with RLE-encoded data provides a computationally efficient modeling alternative to AR1 with uncompressed data.
@@
A related model to the previous one is an ON–OFF survival model, which treats a binary time series as a sequence of state episodes (runs) and models the time spent in each state as a survival outcome. Each run begins when the process enters the ON or OFF state and ends when it switches to the other state. The survival time is the duration in the current state, and the hazard represents the instantaneous risk of switching given how long the state has lasted. By stratifying on transition type (ON→OFF vs OFF→ON), the model allows different baseline hazards for the two states. Covariates can modify these hazards to lengthen or shorten ON or OFF durations. When the time axis is reset at each state entry, the model is semi-Markov, focusing on duration dependence rather than absolute trial time; absolute time can be added as a covariate if needed to capture learning or fatigue. Experimental coefficients describe how the manipulation lengthens or shortens the duration of ON (or OFF) states, not how often states occur or when in the trial switches happen.
@@ 7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmtoolbox [Vanegas2023] was used for analysis of the AR1 model, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 model produced positive definite $R$ matrices; however, the determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ with a ridge of 1e-5. For AR1, estimates were nearly identical to those produced in runs when $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. The data for the analysis were modified in two ways. First, models using a lag-1 predictor required dropping the first observation in each individual time series. Second, for the survival models, the last run was deleted because the transition indicator was missing by fiat. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, COX, and ON-OFF models.
@@
Privileged 0.076 0.071 0.083 0.070 0.069
Contrast -0.079 -0.068 -0.013 -0.059 -0.077
Time 0.854 0.831 0.498 1.816
Priv*Time -0.012 0.006 0.012
Contrast*Time -0.015 -0.012 0.028
  • Note.The 0-to-1 transition effects are reported for the survival models.
@@ 7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the survival analyses, the transition interaction ($y=0\rightarrow y=1)$ effects are given. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model. The GLM, AR1, and ON-OFF models indicate a strong positive relationship between time and the likelihood of being on-target, but the lag-1 $Time$ coefficient is an order of magnitude higher. The variable $Time$ is implicitly contained in the COX model through run start and stop times. For this reason, the COX method did not produce intercepts or time interaction.
It is interesting that the GLM, AR1 and COX/ON-OFF models resulted in similar estimates of experimental effects as the GLM model. Whereas the GLM and AR1 procedures estimate the probability of a on-target gaze, the survival models estimate the transition from on off-target gaze to an on-target gaze.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM and LAG methods provided the lowest, and the AR1 and survival SEs were about double of those for the GLM model. However, the survival models had lower effective sample size because, as noted asbove, the last run in each individual time series is dropped due to the lack of a transition indicator.
@@
Privileged 0.007 0.037 0.014 0.050 0.053
Contrast 0.004 0.023 0.009 0.029 0.031
Time 0.003 0.015 0.006 0.053
Priv*Time 0.007 0.035 0.015
Contrast*Time 0.005 0.022 0.009
  • Note. Robust SEs for survival models are reported for the 0-to-1 transition for the experimental effects.
7.3 Random Effects, $\phi$, and Concordance
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, COX, and ON-OFF models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were moderately similar. For the AR1, the estimated residual correlation was extremely high. However, this is not unexpected given the median run length of 35 (out of 112) across all subjects. This level of $\phi$ raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient. The concordance coefficient was higher for the ON-OFF model (.574 v. .512), meaning the latter model is noticeably better at ranking runs by time-to-switch. Thus, runs that switched earlier were assigned noticeably higher predicted risk (transition to 1) with the ON–OFF model.
Table 4: Random-effects variance, concordance, and associated parameters.
Model Subj Var Item Var Concordance
@@ 7.3 Random Effects, $\phi$, and Concordance
COX 0.512
ON-OFF 0.574
  • Note. Ridge estimates are reported for AR1 with $\phi=.95$. Estimated value of $\phi=.963$ and estimated $\alpha=1.001$. Concordance SEs are both .003.
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM model, the AR1 and survival models provided more about the same estimates of experimental effects. This result raises important and unresolved questions. First, the finding that AR1 and COX/OFF-OFF models provided similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival model [campajola2020ising, gorgi2016survival]. It is an critical caveat that the AR1 model is not truly time series models but rather an error-correlation model. Second, while the autoregressive model with lag coefficients is a useful exploratory tool, it may be impractical for longer time series due to increasing multicollinearity. While a upper boundary can be set manually for the autocorrelation and a ridge added to $|R|$, these are purely ad hoc solutions. The survival models, or similar models, may provide a way to avoid near singularities due to serial correlation. The survival models, with their relationship to Markov models, seem more interesting for further research.
@@
A third issue is that for hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient may have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction (and subject variance component as shown in Table 4).
A fourth issue is the key question of what outcome to model. The GLM, AR1, and LAG models estimate the probability of an on-target gaze, while the survival models focus on the transition from off-target to on-target gazes. This would appear to be a substantive question, and it has important consequences for obtaining standard errors. It would seem less important to model probabilities for adjacent gazes in a run rather than to discover which factors facilitate the transition to an on-target gaze. In this regard, the LAG model initially bears some resemblance to a logistic Markov model, but the resemblance is superficial because a lag-1 indicator does not represent a state transition when binary data are composed of runs. Finally, the strong similarity between the GLM and Survival models for fixed effects may suggest a deeper relationship between modeling on-target gazes and modeling transitions to on-target gazes.
@@
Data pre-processing is another important issue. Eye-tracking systems commonly compress data to avoid creating mammoth output files. Yet even with preliminary machine compression, files may be large. The RLE encoding method for additional compression is convenient for survival analysis and greatly reduces execution time. Pre-processing also includes data cleaning which involves tricky decisions of which experimental records to retain for analysis.
@@
The results of this study collectively support the existence of an experimental effect for the variable $Contrast$. For the AR1 and survival models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength across modeling approaches, this demonstrates model assumptions can have substantively important consequences.
@@ 9 Acknowledgments
I used AI tools in the preparation of this paper. This included preparation of references in biblatex format, review of relevant psychological literature for understanding experimental variables, querying regarding modeling issues, suggesting R code for data file preparation, and obtaining latex code for tables and equations. I generated three paragraphs of text which were heavily edited prior to finalizing the paper. I also used AI to revise a handful of individual sentences.